OpenStudy (anonymous):

Find the slope and y-intercept of y=x+10

OpenStudy (anonymous):

you know the slope-intercept form?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

y = mx + b mx = slope b = intercept

OpenStudy (anonymous):

oh yah i know it but from what i understand to find x and y you substitute 0... so idk if thats right

OpenStudy (mayankdevnani):

i give you some hint

OpenStudy (mayankdevnani):

@monica123

OpenStudy (anonymous):

to find where the cross the axis yes. If you plug 0 = x you will get b. Which is the y-intercept. Same goes if y=0 for x-intercept

OpenStudy (anonymous):

oh..ok

OpenStudy (anonymous):

but the slope can be found in a few ways. Using the point-slope form is an easy way to find it.

OpenStudy (anonymous):

you could aslo do a derivative, but that is calculus stuff.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

glad I could help. so you got it?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

slope is 1 and slope point form y - y1 = m * (x-x1)